Subshilava:

A) Match the area formula to the correct shape. Kite or Rhombus Parallelogram or Rectangle Square Triangle :: A = bh :: A = 1/2 bh :: A = 1/2 d_{1}*d_{2} :: A = s ^ 2

toga:

- Kite or Rhombus: A = 1/2 d_{1}*d_{2} - Parallelogram or Rectangle: A = bh - Square: A = s ^ 2 - Triangle: A = 1/2 bh

Subshilava:

Tysm